# arXiv:2507.20101v2[quant-ph] 02 Mar 2026

### v2

@@
arXiv:2507.20101v2[quant-ph] 02 Mar 2026
@@
Figure 1: Quantum tunnelling between two coupled wavegiudes based on Schrödinger equation. a, Calculated wavefunctions $\psi_{m}$ and $\psi_{a}$ along $x$ direction. Three conditions are considered here with different $\Delta$. The upper panel shows two transmission modes, the middle panel presents one transmission and evanescent modes, and the lower panel displays two evanescent modes. b, Purple dots denotes and cyan line are experimental determined $v$ and fitted results in Ref. [3], respectively. Red line is our theoretical fitting to the experimental results in Ref. [3].
@@
where $\rho_{a}=|\psi_{a}|^{2}/(|\psi_{a}|^{2}+|\psi_{m}|^{2})$. Thus $v$ can be determined by recording the populations $|\psi_{a}|^{2}$ and $|\psi_{m}|^{2}$ in the auxiliary and main waveguides at different $x$ positions, respectively. Noting that this is not velocity measurement since the direction of motion can not be determined. The authors say that this speed “represents the local speed of particles directly behind the potential step”, but the direction of this speed is still unclear. The second velocity is determined using guiding equation $\vec{v}_{s}=\frac{\hbar}{m}\nabla S$ of BM through “interfere quantum states in the step potential with their spatially mirrored images to detect phase gradients”. Based on the authors’ claim of $v_{s}=0$ for negative kinetic energy case, we believe that this $v_{s}$ should along the $x$ direction (see Supplementary information).
@@
The discrepancy of these two measured speeds happens when $\Delta<-\hbar J_{0}$, where $\Delta=E-V_{0}+\hbar J_{0}$. The observed $v$ is nonzero and is proportional to $\sqrt{\frac{2|\Delta|}{m}}$, while measured $v_{s}$ is very close to zero. Therefore, the authors claim that this experiment raises concern on the correctness of Bohmian mechanics.
Here we reanalyze the physical process in this experiment and demonstrate that it poses no challenge to BM. First, the two measured speeds represent fundamentally distinct physical quantities; comparing them is physically unjustified and cannot serve as evidence favoring the Copenhagen interpretation over BM. Second, we theoretically prove that both the widely-held Copenhagen interpretation and BM yield identical predictions for photon tunnelling in coupled waveguides, thereby resolving the contradiction raised by Ref. [3].
@@
To definitively demonstrate this experiment poses no challenge to BM, we prove BM yields identical photon populations $\rho_{a}$ to the Schrödinger solution (equation 4; full derivation in Supplementary information). First, substituting the wavefunction in equation (11) in SM into the Schrödinger equations yields new quantum Hamilton-Jacobi equations and continuity equations. The continuity equations contain a tunnelling current term $j_{0}=2J_{0}R_{a}R_{m}\text{sin}(S_{a}-S_{m})$, which accounts for inter-waveguide quantum tunnelling while preserving probability conservation. Solving the reduced continuity equation $\partial_{x}(\rho_{a,B}v_{a,x})+j_{0}=0$ with BM-derived $v_{a,x}$ and $j_{0}$ gives $\rho_{a,B}$, which is presented in equation (18) in Supplementary information. Subsequent analysis confirms exact agreement with the Schrödinger-predicted $\rho_{a}$ (equation 4). This result indicates that (i) the Bohmian velocity $v_{a,x}$ generates the same $\rho_{a}$ as standard Schrödinger equation, and (ii) both interpretations predict identical tunnelling dynamics, refuting claims in Ref. [3].
In conclusion, the experiment in Ref. [3] poses no challenge to BM. The mistakes in Ref. [3] primarily stem from misinterpreting the semi-classical speed $v$ (equation 2) and Bohmian velocity $v_{s}$ as comparable quantities. We point out that these represent fundamentally distinct physical entities, their discrepancy cannot validate claims against BM. Furthermore, solving for $\rho_{a}$ through both Schrödinger equation and continuity equation in BM yields identical results. This confirms no interpretational distinction exists between Copenhagen and BM for quantum tunneling phenomena. In other words, any velocity defined from $\rho_{a}$ must be identical in both frameworks, meanwhile satisfy continuity equation. Last but not least, $v_{s}=0$ in the region of $\Delta<-\hbar J_{0}$ is physically expected. From the perspective of BM, photons are indeed at rest in the step potential along $x$ direction, but by no means it imply that quantum tunnelling does not occur. In fact, non-vanishing probability of evanescent wavefunction still allows quantum tunnelling to come up.
@@
Tunnelling photons pose no challenge to Bohmian machanics

### v3

@@
Abstract
Very recently, Sharoglazova et al. performed an experiment measuring the energy-velocity relationship and Bohmian velocity in coupled waveguides. Their data show a discrepancy between the semi-classical ‘speed’ $v=\sqrt{2|\Delta|/m}$ and Bohmian velocity $\vec{v}_{s}$ for $\Delta<-\hbar J_{0}$, leading them to claim a challenge to Bohmian mechanics. Here, we definitively demonstrate this experiment poses no challenge to Bohmian mechanics. First, $v$ and $\vec{v}_{s}$ represent fundamentally distinct physical quantities—comparing a scalar and a vector is physically unjustified and cannot adjudicate between Copenhagen and Bohmian interpretations. Second, we rigorously show that both interpretations predict identical photon tunneling dynamics in coupled waveguides.
@@ Abstract
Figure 1: Quantum tunneling between two coupled wavegiudes based on Schrödinger equation. a, Calculated wavefunctions $\psi_{m}$ and $\psi_{a}$ along $x$ direction. Three conditions are considered here with different $\Delta$. The upper panel shows two transmission modes, the middle panel presents one transmission and evanescent modes, and the lower panel displays two evanescent modes. b, Purple dots denotes and cyan line are experimental determined $v$ and fitted results in Ref. [3], respectively. Red line is our theoretical fitting to the experimental results in Ref. [3].
@@
where $\rho_{a}=|\psi_{a}|^{2}/(|\psi_{a}|^{2}+|\psi_{m}|^{2})$. Thus $v$ can be determined by recording the populations $|\psi_{a}|^{2}$ and $|\psi_{m}|^{2}$ in the auxiliary and main waveguides at different $x$ positions, respectively. Noting that this is not velocity measurement since the direction of motion can not be determined. The authors say that this speed “represents the local speed of particles directly behind the potential step”, but the direction of this speed is still unclear. The second velocity is determined using guiding equation $\vec{v}_{s}=\frac{\hbar}{m}\nabla S$ of BM through “interfere quantum states in the step potential with their spatially mirrored images to detect phase gradients”. Based on the authors’ claim of $\vec{v}_{s}=0$ for negative kinetic energy case, we believe that this $\vec{v}_{s}$ should along the $x$ direction (see Supplementary information).
@@
The discrepancy of these two measured speeds happens when $\Delta<-\hbar J_{0}$, where $\Delta=E-V_{0}+\hbar J_{0}$. The observed $v$ is nonzero and is proportional to $\sqrt{\frac{2|\Delta|}{m}}$, while measured $\vec{v}_{s}$ is very close to zero. Therefore, the authors claim that this experiment raises concern on the correctness of BM.
Here we reanalyze the physical process in this experiment and demonstrate that it poses no challenge to BM. First, the two measured speeds represent fundamentally distinct physical quantities: $v$ is a scalar, which quantifies the magnitudes of velocity while the direction is unclear; $\vec{v}_{s}$ is the velocity defined by the continuity equation associated with Schrödinger equation, thus it’s independent with the Copenhagen interpretation or the BM. Comparing a scalar and a vector is physically unjustified and cannot serve as evidence favoring the Copenhagen interpretation over BM. Second, we theoretically prove that both the widely-held Copenhagen interpretation and BM yield identical predictions for photon tunneling in coupled waveguides, thereby resolving the contradiction raised by Ref. [3].
@@
To definitively demonstrate this experiment poses no challenge to BM, we prove BM yields identical photon populations $\rho_{a}$ to the Schrödinger solution (equation 4; full derivation in Supplementary information). First, substituting the wavefunction in equation (11) in SM into the Schrödinger equations yields new quantum Hamilton-Jacobi equations and continuity equations. The continuity equations contain a tunneling current term $j_{0}=2J_{0}R_{a}R_{m}\text{sin}(S_{a}-S_{m})$, which accounts for inter-waveguide quantum tunneling while preserving probability conservation. Solving the reduced continuity equation $\partial_{x}(\rho_{a,B}v_{a,x})+j_{0}=0$ with BM-derived $v_{a,x}$ and $j_{0}$ gives $\rho_{a,B}$, which is presented in equation (18) in Supplementary information. Subsequent analysis confirms exact agreement with the Schrödinger-predicted $\rho_{a}$ (equation 4). This result indicates that (i) the Bohmian velocity $v_{a,x}$ generates the same $\rho_{a}$ as standard Schrödinger equation, and (ii) both interpretations predict identical tunneling dynamics, refuting claims in Ref. [3].
In conclusion, the experiment in Ref. [3] poses no challenge to BM. The mistakes in Ref. [3] primarily stem from misinterpreting the semi-classical speed $v$ (equation 2) and Bohmian velocity $v_{s}$ as comparable quantities. We point out that these represent fundamentally distinct physical entities, their discrepancy cannot validate claims against BM. Furthermore, solving for $\rho_{a}$ through both Schrödinger equation and continuity equation in BM yields identical results. This confirms no interpretational distinction exists between Copenhagen and BM for quantum tunneling phenomena. In other words, any velocity defined from $\rho_{a}$ must be identical in both frameworks, meanwhile satisfy continuity equation. Last but not least, $v_{s}=0$ in the region of $\Delta<-\hbar J_{0}$ is physically expected. From the perspective of BM, photons are indeed at rest in the step potential along $x$ direction, but by no means it imply that quantum tunneling does not occur. In fact, non-vanishing probability of evanescent wavefunction still allows quantum tunneling to come up.
@@
Tunneling photons pose no challenge to Bohmian mechanics